# Training with Pseudo-Code for Instruction Following

### v1

@@ Abstract
Despite the rapid progress in the capabilities of Large Language Models (LLMs), they continue to have difficulty following relatively simple, unambiguous instructions, especially when compositions are involved. In this paper, we take inspiration from recent work that suggests that models may follow instructions better when they are expressed in pseudo-code  (Mishra et al., 2023; Puerto et al., 2024). However, writing pseudo-code programs can be tedious and using few-shot demonstrations to craft code representations for use in inference can be unnatural for non-expert users of LLMs. To overcome these limitations, we propose fine-tuning LLMs with instruction-tuning data that additionally includes instructions re-expressed in pseudo-code along with the final response. We evaluate models trained using our method on $1111$ publicly available benchmarks comprising of tasks related to instruction-following, mathematics, and common-sense reasoning. We conduct rigorous experiments with $55$ different models and find that not only do models follow instructions better when trained with pseudo-code, they also retain their capabilities on the other tasks related to mathematical and common sense reasoning. Specifically, we observe a relative gain of 333319191919% on instruction-following benchmark, and an average gain of upto 14% across all tasks.
**footnotetext: Authors contributed equally to this work.
@@ 1 Introduction
Despite the rapid progress in the capabilities of Large Language Models (LLMs), they continue to struggle following relatively simple instructions, particularly when compositions are involved  (Wen et al., 2024; Jiang et al., 2024b; Murthy et al., 2025). Moreover, they are prone to distractions  (Hwang et al., 2025) and often have difficulty adhering to output format requests  (Xia et al., 2024). To address these challenges, a significant amount of research has focused on improving the instruction-following abilities of LLMs. These include enforcing constraints during decoding  (Banerjee et al., 2025), prompting strategies that elicit thinking and planning behavior through clever design and structure (Xia et al., 2025), test-time scaling (Snell et al., 2024; Bansal et al., 2025), multi-step agentic flows (Yao et al., 2023b; Mitra et al., 2024) and more. Additionally, training time interventions to improve instruction-following is also an active area of work - this includes designing training objectives  (Pham et al., 2024; Shao et al., 2024; Dong et al., 2025; DeepSeek-AI et al., 2025), the inclusion of chain-of-thought  (Longpre et al., 2023b; Mitra et al., 2024), and perhaps, most importantly, curation of training data  (Lou et al., 2024; Sun et al., 2024; Qi et al., 2024; He et al., 2024).
In this paper, we draw inspiration from recent work that suggests models perform better when given instructions in pseudo-code  (Mishra et al., 2023; Li et al., 2024a; Puerto et al., 2024). Pseudo-code instructions offer several advantages over natural language, including being better structured, clearer, and more concise. This clarity and conciseness can improve model performance by reducing the complexity of instructions. Thus, instead of using verbose natural language instructions, prompting models with pseudo-code  (Mishra et al., 2023) or using pseudo-code as a thinking aid during inference  (Puerto et al., 2024) can improve model performance on a wide range of traditional natural-language processing tasks including text classification, question-generation, question-answering, common-sense reasoning, etc.
However, writing pseudo-code programs can be tedious and using few-shot demonstrations to craft code representations for use in inference can be unnatural for non-expert users of LLMs. Additionally, methods such as Chain-of-Code  (Li et al., 2024a) also require the use of external code-emulators for creating line-by-line executed, intermediate outputs of simulated or executed code statements, possibly limiting their adoption.
To address the limitations of such inference-time methods, we fine-tune base LLMs using instruction-tuning data that includes pseudo-code along with the final answer. Thus, given a natural language prompt, models are trained to first re-express the instruction in pseudo-code and then attempt to answer the question (see sample pseudo-code program in Figure 1). From an end-user’s perspective, interacting with such an LLM is identical to regular natural-language prompting, offering a seamless and easy drop-in alternative to existing NL-prompted LMs.
We note that training models to re-express instructions in pseudo-code before answering a user’s request, is similar in spirit, to training models to generate hidden chain-of-thought (CoT) or ‘thinking’ steps (Hayati et al., 2025). However, in this work, we do not enrich pseudo-code programs with such ‘thinking’ steps, and only re-express instructions to be able to isolate the impact of pseudo-code instruction training on LLM performance.
Experimenting with publicly available benchmarks, we find that not only do models follow instructions better when trained with pseudo-code, they also retain and sometimes improve their capabilities on other tasks related to mathematical and common-sense reasoning. Specifically, we observe a relative gain of 333319191919% on Instruction following benchmark, and an average gain of upto 14% across all tasks.
@@
Contributions: In summary, our contributions in this work are as follows: (1) We develop a novel multi-stage data generation pipeline that augments instruction tuning datasets with instructions re-expressed in pseudo-code, (2) We use 0.250.250.250.25M samples from the TuluV2 data mixture (Ivison et al., 2023) to train $55$ different base models. We present comprehensive experiments on $1111$ publicly available benchmarks spanning a range of tasks including instruction-following, common-sense reasoning tasks, and mathematical tasks, (3) We demonstrate that models trained with pseudo-code instructions are better instruction-followers than their counterparts trained exclusively on natural-language instructions, and retain their capabilities on other tasks.
@@ 2 Data Construction
Data Generation Pipeline: As mentioned previously, we enhance existing instruction-tuning datasets by re-expressing instructions in pseudo-code as part of the output. Since manually crafting pseudo-code programs for instruction-tuning is prohibitively expensive, we develop a multi-step data conversion pipeline to create pseudo-code instruction tuning data. Our pipeline consists of three stages: (i) Generate: Use a powerful LLM with a 1-shot prompt to generate pseudo-code programs for a given instruction. We manually write 5 pseudo-code instructions for each of the dataset. We randomly select $11$ of them for the 1-shot prompt. (ii) Evaluate: Similar to Mishra et al. (2023), use only the generated pseudo-code program to run inference and get the output of the instruction. We compare and score the generated output against the ground-truth output, (iii) Repair: For the cases when the inference using the pseudo-code program does not result in the correct ground-truth instruction output, but it does when inference is run using the original natural language instruction, we prompt an LLM-judge to give feedback to improve the pseudo-code program. We summarize the pipeline in Figure 1.
In order to generate pseudo-code programs, the pipeline uses hand-crafted demonstrations of pseudo-code instructions from Mishra et al. (2023) with Mixtral-8x7B-Instruct-v0.1***https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 as the teacher model. Prompts for different stages of this pipeline are available in Appendix Figures 7 and 19. All our pseudo code programs also include a statement mimicking the “invocation” of the program (See Line $1414$ in the Pseudo-code program in Figure 1). An instruction tuning instance consists of the original NL-instruction as the input along with the generated pseudo-code program and the original NL-instruction output text (separated by custom [PSEUDOCODE] markers), as the output of the instruction instance (see example instances in Appendix Section B).
@@
Dataset Source Total #Samples Sampling Proportion
Code Alpaca Chaudhary (2023) 20016 0.15
Tulu v2 Science Rafailov et al. (2024) 7468 0.0375
SNI v2 Wang et al. (2022) 85187 0.325
WizardLM Evol Instruct Xu et al. (2024) 26180 0.15
GPT4-Alpaca Peng et al. (2023) 46149 0.15
CoT Longpre et al. (2023a) 98315 0.1875
Training Mixture: For our experiments, we use datasets from the Tulu v2 mixture (Ivison et al., 2023). We obtain the full set of instances for Code Alpaca (Chaudhary, 2023) and GPT4-Alpaca (Peng et al., 2023) instead of the subset of instances present in Tuluv2 mixture. We choose all instances of Science subset (Rafailov et al., 2024) and WizardLM Evol-Instruct (Xu et al., 2024) from the Tuluv2 mixture. Additionally we use all tasks from SuperNaturalInstructions (Wang et al., 2022) and sample a maximum of $100100$ instances from each task. Lastly, we use all the instances from the CoT subset of the FLAN collection Longpre et al. (2023a). Table 1 lists the datasets used in our experiments along with their statistics.
@@ 2.1 Pseudo-Code Generation
Super-Natural Instructions (SNI v2) dataset: Wang et al. (2022): This is a collection of around $16001600$ traditional NLP tasks formatted as input and output with prompts for instruction-tuning LLMs. We select $55$ pseudo-code demonstrations randomly from Mishra et al. (2023)’s collection of $132132$ tasks for the Generate step. As mentioned in Section 2 for the Evaluate and Repair steps, we prompt the Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2023) model with both pseudo-code instructions and natural languages instructions to generate the answers. We apply additional postprocessing such as splitting based on new line characters etc, to obtain the answer. We then use the exact-match metric to score outputs of the pseudo-code and natural language instructions and trigger the Repair step as needed.
@@
GPT4-Alpaca, Code Alpaca WizardLM Evol-Instruct: For these datasets, we run the pseudo-code generation pipeline without Repair. These datasets usually have open-ended outputs and it is difficult to verify the correctness of the generated response unless we use a robust LLM-as-a-judge. Further, for these instances, forcing updates to the pseudo-code program using the Evaluate and Repair stages runs the risk of leaking output information into the pseudo-code program that can result in “CoT” like steps being included in the program.***This isn’t a problem with SNI v2 because each task has many instances, and instance specific Repairs don’t generalize on other instances. While this isn’t detrimental and in fact likely to be helpful, we wanted to be careful and ensure we are able to study the impact of training with NL-equivalent pseudo-code instructions without any enrichment. Therefore, we only use the Generate part of our pipeline for these datasets. We select $55$ instances from each dataset and write the pseudo-code demonstrations ourselves. The instances are chosen to be diverse; for the the $11$-shot prompt, we randomly select one of these pseudo-code demonstrations.
@@
TuluV2 Science and CoT Data: We manually write pseudocode instructions for each of the $99$ datasets present in the CoT collection Longpre et al. (2023a) and the Tulu V2 Science subsets Ivison et al. (2023). The pseudo-code programs for each of these datasets are provided in the Appendix.
3 Experiments
Through our experiments we seek to answer the following research questions: (i) Does instruction-tuning with pseudo-code help improve the performance of models on instruction-following tasks?, (ii) Does the introduction of pseudo-code during instruction-tuning affect the ability to acquire skills imparted during traditional instruction-tuning, (iii) The nature of our pseudo-code is very python-like and we also do not use a lot of code data in our training mixture. To what extent does this affect the performance on code-tasks?
3.1 Models
Since we study the effect of instruction-tuning with and without pseudo-code instructions, we require unaligned base-models. After a careful review of available choices, we settled on the 78787-87 - 8B parameter range models. Additionally, prior work by Mishra et al. (2023) suggests that medium parameter-size models (eg: 6B-7B) were able to exploit pseudo-code better than smaller models. We experiment with the following $55$ base models which includes $11$ code base model: (i) meta-llama/Llama-3.1-8B, ***https://ai.meta.com/blog/meta-llama-3-1/ (ii) mistralai/Mistral-7B-v0.3,***https://huggingface.co/mistralai/Mistral-7B-v0.3 (iii) Qwen/Qwen2.5-7B  (Qwen et al., 2025), (iv) ibm-granite/granite-3.1-8b-base***https://huggingface.co/ibm-granite/granite-3.1-8b-base and (v) ibm-granite/granite-8b-code-base-128k (Mishra et al., 2024).
Training: All models were SFT trained on a full NVIDIA A100 80GB node with Flash Attention $22$ (Dao, 2024) and bf16 precision using the Dolomite Engine Framework.***https://github.com/IBM/dolomite-engine All models were trained for $22$ epochs using the PyTorch’s implementation of AdamW  (Loshchilov and Hutter, 2019) as the optimizer, with a starting learning rate of 1e-5 and linear decay. Models were trained with two parallel training mixtures - one with NL instructions and the other augmented with pseudo-code programs.
3.2 Evaluation Tasks
We conduct experiments on three main types of tasks (i) Instruction-Following, (ii) Common sense and Reasoning, (iii) Mathematical Tasks.
Instruction Following: Since our primary goal was to investigate if instruction-tuning with pseudo-code helps instruction-following, we conduct extensive experiments on instruction-following benchmarks including IFEval  (Zhou et al., 2023), KCIF  (Murthy et al., 2025), Complex Bench  (Wen et al., 2024) and Follow Bench  (Jiang et al., 2024b).
IFEval and KCIF both use non-LLM based scoring functions to evaluate performance on instruction-following; ComplexBench  (Wen et al., 2024) and FollowBench (Jiang et al., 2024b) use a powerful LLM-as-a-judge to assess instruction-following. While Complex bench helps study different instruction compositions (eg: conjunctions, chains, conditionals, etc), FollowBench reports performance aggregated by task categories (eg: instruction constraints applied to response styles, situations, etc). Further, IFEval only assesses whether the lexical and formatting instructions are followed in responses with no consideration of task completion. All other datasets condition instruction following on core-task performance - eg: KCIF (Murthy et al., 2025) has answer modifying instructions from knowledge tasks, where models are penalized if instructions are applied on the wrong answer to a question.
Common Sense & Reasoning Tasks: We also study the performance of models on commonly used benchmarks to investigate whether our method interferes with any skills acquired by vanilla NL-SFT. We present experiments on Reasoning tasks (Hellaswag  (Zellers et al., 2019), PiQA  (Bisk et al., 2020), Winogrande  (Sakaguchi et al., 2021) and ARC  (Clark et al., 2018)), and Mathematical Tasks (GSM8K  (Cobbe et al., 2021), as well as a revised version of GSM8K called GSM8K Platinum  (Vendrow et al., 2025), and MathQA  (Amini et al., 2019)).
Metrics: We post-process outputs and then use the official scoring scripts from the benchmarks where available. IFEval has custom-scorers for each instruction and it aggregates results at the prompt as well as instruction level. ComplexBench and FollowBench aggregate LLM-judge generated scores grouped by different task categories. KCIF as well as, the Common Sense and Reasoning tasks are measured in terms of exact-match scores.
3.2.1 Evaluation Setup and Metrics
Output Post Processing: We develop robust task-specific output-processing pipelines for both NL and pseudo-code models. (See Appendix Section F for more details). All tasks are evaluated after output post-processing.
3.3 Results
@@
Table 2: Comparison of NL-tuned vs pseudo-code tuned models on Instruction-Following, Mathematics, Commonsense & Reasoning and Knowledge Tasks.
Instruction-Following Mathematics
Common Sense
& Reasoning

### v2

@@ Abstract
Despite rapid advances in the capabilities of Large Language Models (LLMs), they continue to struggle with following relatively simple and unambiguous instructions, particularly when compositional structure is involved. Recent work suggests that models may follow instructions more effectively when they are expressed in pseudo-code rather than natural language (Mishra et al., 2023; Puerto et al., 2024). However, writing pseudo-code programs can be tedious, and relying on few-shot demonstrations or inference-time code prompting is often unnatural for non-expert users of LLMs. To overcome these limitations, we propose a training-time approach that fine-tunes LLMs using instruction-tuning data augmented with pseudo-code representations of natural-language instructions paired with final responses. We evaluate our method on 12 publicly available benchmarks spanning instruction-following, mathematical reasoning, and commonsense reasoning, across six base models. Our results show that models trained with pseudo-code follow instructions more reliably, achieving relative gains of 8–21% on instruction-following benchmarks, while largely preserving—and in some cases improving—performance on mathematical and commonsense reasoning tasks, with an average gain of up to 30% across all evaluated benchmarks.
Machine Learning, ICML
@@ 1 Introduction
Despite the rapid progress in the capabilities of Large Language Models (LLMs), they continue to face challenges in effectively processing and executing instructions across diverse tasks. These challenges manifest in several ways: difficulty with compositional instructions (Wen et al., 2024; Jiang et al., 2024b; Murthy et al., 2025), susceptibility to distractions (Hwang et al., 2025), and struggles with adhering to specified output formats (Xia et al., 2024). To address these limitations, significant research effort has focused on improving how models understand and respond to instructions. This includes enforcing constraints during decoding (Banerjee et al., 2025), prompting strategies that elicit structured thinking and planning behavior (Xia et al., 2025), test-time scaling (Snell et al., 2024; Bansal et al., 2025), multi-step agentic flows (Yao et al., 2023b; Mitra et al., 2024), and more. Additionally, training-time interventions have emerged as an active area of work, including the design of training objectives (Pham et al., 2024; Shao et al., 2024; Dong et al., 2025; DeepSeek-AI et al., 2025), incorporation of chain-of-thought reasoning (Longpre et al., 2023a; Mitra et al., 2024), and, crucially, careful curation of training data (Lou et al., 2024; Sun et al., 2024; Qi et al., 2024; He et al., 2024).
@@
In this paper, we draw inspiration from recent work showing that large language models often perform better when instructions are expressed in pseudo-code rather than natural language (Mishra et al., 2023; Li et al., 2024a; Puerto et al., 2024). Compared to natural language, pseudo-code provides a more explicit representation of task structure, including ordering, control flow, and execution constraints. Mishra et al. (2023) further show that code-specific elements—such as comments, docstrings, and structural cues, can improve model performance. By clarifying task structure and constraints, these elements yield instructions that are clearer, more concise, and less ambiguous. As a result, both directly prompting models with pseudo-code (Mishra et al., 2023) and using pseudo-code as a thinking aid during inference (Puerto et al., 2024) have been shown to improve performance across a wide range of natural language processing tasks, including text classification, question answering, question generation, and commonsense reasoning.
However, relying on pseudo-code at inference time introduces practical limitations. Writing pseudo-code programs can be tedious and unintuitive for non-expert users, and few-shot demonstrations that incorporate code representations are often brittle. Moreover, approaches such as Chain-of-Code (Li et al., 2024a) depend on external code execution or emulation to produce intermediate outputs, complicating deployment.
To address these limitations, we propose a training-time alternative. We fine-tune base LLMs using instruction-tuning data that pairs natural-language instructions with corresponding pseudo-code representations and final answers. Given a natural-language prompt, the model is trained to first re-express the instruction in pseudo-code and then generate the answer (see the sample program in Figure 1). From an end-user’s perspective, interaction with the resulting model is identical to standard natural-language prompting, making our approach a seamless drop-in replacement for existing instruction-tuned LLMs.
Training models to re-express instructions in pseudo-code before answering a user’s request is similar in spirit to approaches that train models to generate hidden chain-of-thought (CoT) or “thinking” steps (Hayati et al., 2025). However, in this work we do not enrich pseudo-code instructions (programs) with such “thinking” steps; instead, we re-express instructions solely to isolate the impact of pseudo-code–based instruction training on LLM performance.  [^1] Note: Our training data includes some tasks where CoT is expressed in final responses (but not in programs).
Through experiments on publicly available benchmarks, we find that models trained with pseudo-code–augmented instruction tuning follow instructions more reliably and robustly than their natural-language–trained counterparts. Moreover, these gains do not come at the expense of general capability: pseudo-code–trained models retain—and in some cases improve—their performance on tasks other tasks such as those involving mathematical and commonsense reasoning.
Contributions: We propose a training-time approach for improving instruction-following in large language models by learning a structured pseudo-code intermediate representation of natural-language instructions. Specifically, (1) we train models to re-express instructions in pseudo-code before answering, improving execution of compositional, constrained, and nested instructions without inference-time prompting or external code execution; (2) we introduce a scalable generate–evaluate–repair pipeline that automatically augments instruction-tuning data with high-quality pseudo-code supervision; and (3) through experiments on six base models and twelve benchmarks, we show performance improvements of (8–21% relative) on instruction-following tasks while largely preserving or improving performance on mathematical, commonsense, and code-related tasks, and outperforming inference-time code prompting baselines.
@@ 2 Data Construction
Data Generation Pipeline: As mentioned previously, we enhance existing instruction-tuning datasets by re-expressing instructions in pseudo-code as part of the output. Since manually crafting pseudo-code programs for instruction-tuning is prohibitively expensive, we develop a multi-step data conversion pipeline to create pseudo-code instruction tuning data. Our pipeline consists of three stages: (i) Generate: Use a powerful LLM with a 1-shot prompt to generate pseudo-code programs for a given instruction. We manually write 5 pseudo-code instructions for each of the datasets. We randomly select $1$ of them for the 1-shot prompt. (ii) Evaluate: Similar to (Mishra et al., 2023), use only the generated pseudo-code program to run inference and get the output of the instruction. We compare and score the generated output against the ground-truth output, (iii) Repair: For the cases when the inference using the pseudo-code program does not result in the correct ground-truth instruction output, but it does when inference is run using the original natural language instruction, we prompt an LLM-judge to give feedback to improve the pseudo-code program. We summarize the pipeline in Figure 1.
In order to generate pseudo-code programs, the pipeline uses hand-crafted demonstrations of pseudo-code instructions from (Mishra et al., 2023) with Mixtral-8x7B-Instruct-v0.1222https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 as the teacher model. Prompts for different stages of this pipeline are available in Appendix Figures 7 and 30. All our pseudo code programs also include a statement mimicking the “invocation” of the program (See Line $13$ in the Pseudo-code program in Figure 1). An instruction tuning instance consists of the original NL-instruction as the input along with the generated pseudo-code program and the original NL-instruction output text (separated by custom [PSEUDOCODE] markers), as the output of the instruction instance (see example instances in Appendix Section D).
Dataset #Samples Sampling Proportion Code Alpaca (Chaudhary, 2023) 20016 0.15 Tulu v2 Science (Rafailov et al., 2024) 7468 0.0375 SNI v2 (Wang et al., 2022) 85187 0.325 WizardLM Evol Instruct (Xu et al., 2024) 26180 0.15 GPT4-Alpaca (Peng et al., 2023) 46149 0.15 CoT (Longpre et al., 2023b) 98315 0.1875
@@
Training Mixture: For our experiments, we use datasets from the Tulu v2 mixture (Ivison et al., 2023). We obtain the full set of instances for Code Alpaca (Chaudhary, 2023) and GPT4-Alpaca (Peng et al., 2023) instead of the subset of instances present in Tuluv2 mixture. We choose all instances of Science subset (Rafailov et al., 2024) and WizardLM Evol-Instruct (Xu et al., 2024) from the Tuluv2 mixture. Additionally we use all tasks from SuperNaturalInstructions (Wang et al., 2022) and sample a maximum of $100$ instances from each task. Lastly, we use all the instances from the CoT subset of the FLAN collection (Longpre et al., 2023b). Table 1 lists the datasets used in our experiments along with their statistics.
@@ 2.1 Pseudo-Code Generation
Super-Natural Instructions (SNI v2) dataset: (Wang et al., 2022): This is a collection of around $1600$ traditional NLP tasks formatted as input and output with prompts for instruction-tuning LLMs. We select $5$ pseudo-code demonstrations randomly from (Mishra et al., 2023)’s collection of $132$ tasks for the Generate step. As mentioned in Section 2 for the Evaluate and Repair steps, we prompt the Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2023) model with both pseudo-code instructions and natural languages instructions to generate the answers. We apply additional postprocessing such as splitting based on new line characters etc, to obtain the answer. We then use the exact-match metric to score outputs of the pseudo-code and natural language instructions and trigger the Repair step as needed.
@@
GPT4-Alpaca, Code Alpaca WizardLM Evol-Instruct: For these datasets, we run the pseudo-code generation pipeline without Repair. These datasets usually have open-ended outputs and it is difficult to verify the correctness of the generated response unless we use a robust LLM-as-a-judge. Further, for these instances, forcing updates to the pseudo-code program using the Evaluate and Repair stages runs the risk of leaking output information into the pseudo-code program that can result in “CoT” like steps being included in the program. [^3] This isn’t a problem with SNI v2 because each task has many instances, and instance specific Repairs don’t generalize on other instances. While this isn’t detrimental and in fact likely to be helpful, we wanted to be careful and ensure we are able to study the impact of training with NL-equivalent pseudo-code instructions without any enrichment. Therefore, we only use the Generate part of our pipeline for these datasets. We select $5$ instances from each dataset and write the pseudo-code demonstrations ourselves. The instances are chosen to be diverse; for the the $1$-shot prompt, we randomly select one of these pseudo-code demonstrations.
@@
TuluV2 Science and CoT Data: We manually write pseudocode instructions for each of the $9$ datasets present in the CoT collection (Longpre et al., 2023b) and the Tulu V2 Science subsets (Ivison et al., 2023). The pseudo-code programs for each of these datasets are provided in the Appendix.
The data loop illustrated in Figure 1 follows an iterative generate–evaluate–repair paradigm for improving instruction-following behavior. Starting from a natural language (NL) instruction, the system generates a corresponding pseudo-code program using one-shot prompting. Both the NL instruction and the generated pseudo-code are independently evaluated via LLM inference and a scorer, yielding scores $S_{\text{NL}}$ and $S_{\text{PC}}$. When the pseudo-code underperforms (i.e., $S_{\text{NL}}$$S_{\text{PC}}$), targeted feedback is applied to repair and refine the program, thereby closing the loop. Only instruction-pseudo-code pairs that consistently satisfy the evaluation criteria are retained and stored along with their ground-truth outputs.
The effectiveness of this data loop is reflected in the evolution of dataset statistics of SNI v2 across iterations. The initial pass (Iteration 0) contains $1613$ instances generated with minimal filtering. After one full loop of evaluation and repair, Iteration 1 reduces the set to $662$ instances, indicating substantial pruning of low-quality or misaligned samples. A subsequent iteration further refines the dataset to $531$ instances. This steady reduction demonstrates that successive data loop iterations enforce increasingly strict quality thresholds, resulting in a high-precision dataset optimized for downstream training and evaluation.
@@
Table 2: Comparison of Code-Prompted (CP) (Puerto et al., 2024), NL-tuned and pseudo-code (PC) tuned models on Instruction-Following, Mathematics, Commonsense & Reasoning tasks.
Instruction-Following Mathematics Common Sense & Reasoning Avg